OpenStudy (anonymous):

HELP! PLEASEEEE! WILL FAN + METAL Solve the system by substitution 3x-4y= -5 x= 2y-3

OpenStudy (anonymous):

as you have 3x-4y = -5 here you put 2y-3 in the place of x because x= 2y-3 thus we get to 3(2y-3) -3 =0 now solve for y

OpenStudy (anonymous):

put the vale of x in the second equation in the first equation and then solve for y

OpenStudy (anonymous):

after that pick any equation put the value of y and then solve for x

OpenStudy (anonymous):

3(2y-3) -4y = -5 solve for y

OpenStudy (anonymous):

y=7?